# Exploring the Mechanisms Behind Non Aromatic Fluorescence with Density Functional Tight Binding Method

### 20 Jan 2024 v2

@@ 3.4 QM/MM Simulations
In addition to simulations of the dimer in the gas phase, we also conducted the QM/MM simulations in order to understand the role of the environment64, 65. This method involves partitioning the system into two distinct regions, namely a QM and MM, and describing the coupling between these regions using the electrostatic embedding approach66, 65, 67. To investigate the radiative and non-radiative process in both systems, we selected two different QM sub-systems for L-pyro-amm, the dimer and unit cell, while for L-gln we studied the unit cell.
To construct the system, we began by replicating the unit cell of the crystal structure in all three directions resulting in a supercell comprising 375 unit cells for both L-gln and L-pyro-amm. The subsequent step involved a classical equilibration at room temperature for 10 ns with a timestep of 1 fs. Classical parameters for bonded and Lennard-Jones were adopted from GAFF60 and for treating the electrostatics we used Mulliken charges which were extracted from optimized crystal structures using DFTB in PBC (see section 3.1). A total of 10 snapshots were extracted from the simulations.
After selecting the QM region, QM/MM ground state dynamics simulation of 25 ps each were conducted with the same simulation protocol for each snapshot. From the last 20 ps of these simulations, we extracted 20 frames for each trajectory to calculate the absorption spectra using TD-DFTB in the crystal environment (see section 3.2). A total of 200 snapshots were utilized for the absorption spectra.
In addition, we performed TSH simulations for each of the QM sub-system with the aim to study the emission as well as the photo-physical relaxation of the $S_{1}$ state of these systems. TSH simulations started from the $S_{1}$ electronic state employing the initial conditions generated in the previous step for a total of 1 ps, with a timestep of 0.5 fs within the NVE ensemble. A total of 10 TSH dynamics were conducted for each system using the same methodology described section 3.3.
@@ 4.1 Ground State Properties in Molecular Crystals
Before analyzing the optical properties of these non-aromatic systems with TD-DFTB, it is necessary first to verify the quality of the hydrogen bonds (HBs) description, in particular the difference in the distance and its strength for the L-Pyro-Ammonium (L-pyro-amm) and L-Glutamaine (L-gln) in the crystal structures, since this property plays an important role in the fluorescence12, 9. Figure 1 shows the probability distribution function (PDF) obtained from ab-initio ground state molecular dynamics with PBC (see section 3.1) for the typical descriptors employed to describe the HBs including the proton transfer coordinate (PTC) and bond distance (BD) between the two heavy atoms among which proton transfer potentially occurs. BD is defined as the distance between O and N atoms in the case of L-gln (panel a and c) and the distance between the two O atoms in the case of L-pyro-amm (panel b and d). The PTC is defined as the distance difference between the O-H and N-H in L-gln (panel e) and between the two O-H involved in L-pyro-amm (panel f).
Figure 1: Unit cell crystallographic geometries for L-gln (panel a) and L-pyro-amm (panel b), where the blue circles highlight the HBs and SHBs, respectively. Histograms of BD (panels c and d) and PTC (panels e and f) obtained from an ab-initio ground state dynamics with PBC using DFT (blue line) and DFTB (red lines)
@@ 4.2 Absorption Spectra in Molecular Crystals
With the ground-state configurations obtained from the ab-initio MD, we next turn to extracting the absorption spectra. Figure 2 shows the absorption spectra for L-gln and L-pyro-amm obtained experimentally12 (left panel) as well as the calculated with TD-DFTB (right panel). As one can observe, an important difference between the two systems is the extent of the red-tail/edge excitation that is significantly more pronounced in the case of L-pyro-amm. In the case of L-gln, the absorption peaks are below the 250 nm and outside the window of experimentally detectable wavelengths. On other hand for L-pyro-amm, there is a very pronounced red-edge excitation tail ranging from 250-400 nm.
Figure 2: Absorption UV-Vis Spectra for L-Glutamine and L-Pyro-Ammonium. Left panel shows the experimental data, extracted from reference 12. Right panel shows the calculated average spectra using TD-DFTB. Red lines correspond to L-pyro-amm system while blue lines for L-gln.
The right panel of Figure 2 depicts the average calculated absorption spectra obtained for both systems using 1000 snapshots each, at the level of TD-DFTB. Interestingly, we observe that TD-DFTB reproduces the correct experimental trends comparing L-gln and L-pyro-amm. Analysing the transitions involved in the UV-Vis energy range (330-370 nm) we found that the main molecular orbitals that explain these transitions involve significant contributions coming from the carbonyl bonds and the SHB (see Figure S2 in the Supplementary Information). It is also worth noting that this prediction is also fully consistent with the results that were previously obtained at TD-DFT level in our previous work12.
@@ 4.3 Excited State TSH Simulations in Gas phase
Figure 3: Average populations of the electronic states obtained from TSH trajectories for L-gln (left panel) and L-pyro-amm (right panel) dimer systems in vacuum. Results using TD-DFT are depicted as dotted lines while those obtained with TD-DFTB in solid lines. The evolution of $S_{1}$ and $S_{0}$ are denoted in red and blue lines, respectively.
@@
In our previous studies employing TD-DFT, we observed that the transition from $S_{1}\to S_{0}$ occurs via a conical intersection (CoIn) where the essential vibrational modes coupled to the process involves the extension of the carbonyl bond, proton transfer along the SHB, and finally, the deplanarization of the amide group. Examining each of these modes in the ground and excited state as well as near the CoIn, a very specific structural signature becomes apparent. Figure 4 depicts the histograms for each of these modes obtained from TSH simulations employing both TD-DFT (top) and TD-DFTB (bottom) theory levels for L-gln dimer.
@@
Figure 4: Structural features characterizing ground and excited states as well as CoIn crossing for L-gln system. Molecular structures showing the parameters to describe the carbonyl stretching (panel a), the proton transfer (panel d) and the deplanarization (panel g). Probability Distribution Function (PDF) for the CO distance (panels b and c), PTC (panels e and f) and Deplanarization modes (panels h and i) for L-gln dimers in vacuum obtained from TSH trajectories using TD-DFT (middle panels) and TD-DFTB (lower panels). All PDFs were normalized independently.
@@
The comparison between TD-DFTB and TD-DFT reveals excellent agreement both in terms of the overall trends observed but also in the specific details of how the modes change during non-radiative decay. This is a clear evidence of the high degree of distortions experienced by the system upon excitation, in which the minimum on both potential energy surfaces $S_{1}$ and $S_{0}$ are shifted between each other, a feature that is captured by the Tight Binding (TB) approach. In particular, we observe that an additional stretching for CO respect to $S_{1}$ distribution is necessary for the system to reach the CoIn (green lines in panel b and c). Similar behavior is also observed for the PTC (panels e and f). The primary difference arises in relation to the deplanarization mode. In TD-DFT (panel h), we observe a higher degree of deplanarization at the CoIn compared to the distribution of this mode in the $S_{0}$ state. In contrast, although the distributions obtained with TD-DFTB reflect the same trends (panel i), the deplanarization mode appears to be much more constrained compared to our previous findings. Despite the differences, it should be underscored that, based in our previous work16, the deplanarization mode is less critical for the photophysics of both L-pyro-amm and L-gln.
@@
We next turn to comparing the behavior of the non-radiative decay modes that are important for L-pyro-amm. Since this system is fluorescent, extracting the decay mechanisms that emerge from the excited state simulations is particularly challenging due to the fact that there are fewer statistics of reactive events. In our previous study16 for example, out of a total of 100 trajectories, 4 displayed non-radiative decay from the excited to the ground-state. Employing excited states simulations for L-pyro-amm as well as model peptides inspired from the amyloid proteins, we demonstrated that a key ingredient for enhancing excited state lifetime is facilitated when the position of the minima for the $S_{0}$ and $S_{1}$ states overlap, which effectively makes the conical intersection much less accessible16.
@@
Figure 5: Structural features characterizing ground and excited states as well as CoIn crossing for L-pyro-amm system. Molecular structures showing the parameters to describe the carbonyl stretching and the proton transfer are in panel a and d, respectively. Probability Distribution Function (PDF) for the CO distance (panels b and c), PTC (panels e and f) for L-pyro-amm dimers in vacuum obtained from TSH trajectories using TD-DFT (middle panels) and TD-DFTB (lower panels). All PDFs were normalized independently.
@@
In Figure 5 we show the most relevant features, namely the carbonyl stretching and the proton transfer modes obtained from the TSH simulations at TD-DFT and TD-DFTB levels. Specifically, we are able to reproduce the significant overlap in the distribution of the two modes between the $S_{1}$ and $S_{0}$ electronic states (red and blue lines). These findings indicate that the energy minimum in the potential energy surfaces for each of these modes do not exhibit significant displacement, suggesting that these modes are restricted in L-pyro-amm. This feature is also fully consistent with typical aromatic fluorescent systems68, where the restriction arises from the high electron delocalization in the planar arrangement of these compounds. In the case of non-aromatic systems, like the ones we are tackling, the strong interactions due to SHBs induce this restriction on different modes.
@@
It is important to highlight when comparing the top and bottom panels of Figure 5, that due to the reduced computational complexity, we are able to run significantly more TSH trajectories compared to our previous TDDFT simulations (600 vs 100). DFTB thus permits the observation of many more non-radiative decays yielding statistics which allow the construction of histograms of various modes near the CoIn, for instance the green distribution in the case of TD-DFTB represents 122 decays (panels c and f) and the green vertical lines for TD-DFT represents only 4 decays (panels b and e). The examination of the distribution of these modes at the CoIn reveals that the mechanism observed in L-pyro-amm is similar to that observed in the L-gln system. Specifically, we find with the DFTB approach that the CO stretching mode undergoes two distinct signatures at the CoIn. One exhibits a more subtle stretching in comparison to the distribution of $S_{1}$, while the other shows a more significant stretching of approximately $1.6\AA$ (green line in panel c in Figure 5). Our previous TDDFT simulations sample a few events between these two peaks.
@@
The effect of having more statistics with DFTB is observed more clearly in the PTC distributions (panels e and f) which display more structure compared to those obtained with TD-DFT. For example, we can capture the double-well character of the PTC (blue line in panel f) with TD-DFTB, while with TD-DFT we observe one main peak along with a shoulder (blue line panel e). A distinction emerges when comparing the behavior of proton transfer between the excited and ground states in TD-DFTB. Notably, in the excited state, there is a significant reduction in proton transfer events, evidenced by the presence two peaks located at ±$\pm 0.5$with difference intensity in the PDF (depicted by the red line in panel f). A similar trend is observed in TDDFT, albeit with less intensity. In this case, a narrow distribution is noted in $S_{1}$ compared to the ground state, with the hydrogen positioned equidistant from the two oxygen atoms involved in the SHB (illustrated by the red line in panel e). Additionally, we observe that the SHB experiences a deformation at the CoIn (green line in panel f in Figure 5), transforming into a transient normal HB. This is evident from the peaks at higher values of approximately ±$\pm 2.5\AA$, which is also present in TD-DFT (green bars in panel e). Although we cannot rule out the possibility that some of these differences arise from the underlying potential energy surface along various vibrational modes, the limited statistics with TD-DFT are more likely to be the source of the differences we observe. However, we can see that the mechanism predicted by TD-DFTB is in excellent agreement with our previous works12, 16.
@@ 4.4 Optical Properties in Realistic Biological Environments
The preceding results provide evidence that DFTB is a very viable approach for studying the non-radiative decay mechanisms in organic molecular crystals implicated in non-aromatic fluorescence. In the following section, we turn to using TD-DFTB to study the optical properties of L-gln and L-pyro-amm where the environment is included in a more realistic manner within a QM/MM framework.
4.4.1 L-Pyro-Ammonium
@@
The experimental data used to compare the optical properties in this system was the excitation and emission spectra, as reported in ref 12. Therefore, to allow for a one-to-one comparison, we determined the excitation spectra instead of the absorption. To compute the absorption spectra, one has to include all the electronic states in the UV-Vis range (250-400 nm), while to determine the excitation spectra in general one includes only the first few transitions. We chose to compute the excitation spectra using only the first transition ($S_{0}\to S_{1}$). This decision is grounded in the well-established mirror image rule69, which states that if the emission spectrum of a fluorophore mirrors its absorption or excitation spectra, implying that the probability of the $S_{1}\to S_{0}$ transition matches that of the $S_{0}\to S_{1}$ transition. Indeed we observe that this feature is observed in the experimental spectra shown in the right panel of Figure 7.
@@
Figure 7: Excitation (solid lines) and Fluorescence (dashed lines) spectra for L-pyro-amm. Left panel shows the average spectra obtained from QM/MM simulations using the dimer (red line) and the unit cell (blue line) in the QM region. Right panel shows the experimental spectra (extracted from reference 12, red line) and the average calculated from the QM/MM simulations for the unit cell in the QM region, where the calculated spectra were shifted to match the maximum in the experimental fluorescence spectra.
@@
Left panel in figure 7 shows the results for both QM systems in L-pyro-amm in the QM/MM simulation. We can observe that both systems show similar excitation maximum (approx. 215 nm) and emission (approx. 290 nm). When we compared these values with the experimental data (red lines in the right panel) we observe notable differences. This finding is a well know fact that TD-DFT and therefore in its approximated variant TD-DFTB, tend to exhibit a notable shift in the absolute magnitude of the transition energies when comparing with experimental measurements70, 71. An additional factor contributing to the discrepancy between our computational outcomes and experimental findings lies in the inherent limitations of the QM/MM approach in representing the environmental effects. As demonstrated in Section 4.2, our methodology achieves outstanding agreement when employing ab-initio ground state simulations with periodic boundary conditions. Regrettably, this level of concordance is not replicated in our QM/MM simulations. However, as we will elucidate in the following, these disparities persist not only in the ground state but also in excited state simulations. Consequently, focusing on energy differences becomes a more pertinent and meaningful avenue for comparison. We have thus determined the Stokes Shift, defined as the difference between the maximum of the emission band and the absorption or excitation spectra the results of which are summarized in Table 1. Notably, we find remarkable agreement between the experimental results and the two model systems for the Stokes shift. It is also interesting to observe that the dimer system appears to fully capture the energetics associated with the fluorescence Stokes Shift which reinforces the use of the dimer model systems in disentangling the non-radiative decay mechanisms.
@@
Table 1: Stokes Shifts for the different models in the QM/MM simulations
  • $a$ FWHM denotes the Full Width at Half Maximun for the fluorescence spectra.
@@
The right panel of Figure 7 presents the excitation and fluorescence spectra for both experimental and calculated using the unit cell of L-pyro-amm in the QM/MM simulations. The calculated spectra were adjusted to align with the peak of the emission spectra. Impressively, our calculations are in rather modest agreement with the experimental data. The deviation of 10 nm in the Stokes Shift is attributed to the fact that the mirror image rule isn’t entirely applicable in this context. In particular, slight shoulders are noticeable in the experimental excitation spectra (solid red line), which are absent in the emission (dotted red line). This is an indication that the excitation spectra involves more that only one transition from the ground state. Consequently, we computed the excitation spectra by incorporating an increasing number of transitions to accommodate this behavior. The detailed results are illustrated in Figure S3 and table S1 of the Supplementary Information.

### 09 Apr 2024 v3

@@ 3.4 QM/MM Simulations
The accurate simulation of the photophysical relaxation of the excited states in these crystals require the consideration of PBC. However, DFTB+ code does not calculate forces in the excited states within these conditions. One potential solution to circumvent this challenge is to conduct simulations on clusters in the gas phase, as discussed in the preceding section. Another more effective alternative is to employ QM/MM simulations to elucidate the influence of the environment65, 66. This method involves partitioning the system into two distinct regions, namely a QM and MM, and describing the coupling between these regions using the electrostatic embedding approach67, 66, 68. To investigate the radiative and non-radiative process in both systems, we selected two different QM sub-systems for L-pyro-amm, the dimer and unit cell, while for L-gln we studied the unit cell.
To construct the system, we began by replicating the unit cell of the crystal structure in all three directions resulting in a supercell comprising 375 unit cells for both L-gln and L-pyro-amm. The subsequent step involved a classical equilibration at room temperature for 10 ns with a timestep of 1 fs. Classical parameters for bonded and Lennard-Jones were adopted from GAFF61 and for treating the electrostatics we used Mulliken charges which were extracted from optimized crystal structures using DFTB in PBC (see section 3.1). A total of 10 snapshots were extracted from the simulations.
For each snapshot, we proceed to select the QM region situated at the center of our supercell, then a QM/MM ground state dynamics simulation of 25 ps each were conducted with the same simulation protocol. From the last 20 ps of these simulations, we extracted 20 frames for each trajectory to calculate the absorption spectra using TD-DFTB in the crystal environment (see section 3.2). A total of 200 snapshots were utilized for the absorption spectra.
In addition, we performed TSH simulations for each of the QM sub-system with the aim to study the emission as well as the photo-physical relaxation of the $S_{1}$ state of these systems. TSH simulations started from the $S_{1}$ electronic state employing the initial conditions generated in the previous step for a total of 1 ps, with a timestep of 0.5 fs within the NVE ensemble. A total of 10 TSH dynamics were conducted for each system using the same methodology described section 3.3. For the fluorescence spectra, we employed trajectories that did not show decay to the ground state and excluded the initial 200 fs from the analysis, which corresponds to the time for the system to reach the minimum on $S_{1}$. The emission spectra were computed by utilizing the energy gap between $S_{1}$ and $S_{0}$, along with the oscillator strength, and then averaging.
@@ 4.1 Ground State Properties in Molecular Crystals
Before analyzing the optical properties of these non-aromatic systems with TD-DFTB, it is necessary first to verify the quality of the hydrogen bonds (HBs) description, in particular the difference in the distance and its strength for the L-Pyroglutamine-Ammonium (L-pyro-amm) and L-Glutamaine (L-gln) in the crystal structures, since this property plays an important role in the fluorescence12, 9. Figure 1 shows the probability distribution function (PDF) obtained from ab-initio ground state molecular dynamics with PBC (see section 3.1) for the typical descriptors employed to describe the HBs including the proton transfer coordinate (PTC) and bond distance (BD) between the two heavy atoms among which proton transfer potentially occurs. BD is defined as the distance between O and N atoms in the case of L-gln (panel a and c) and the distance between the two O atoms in the case of L-pyro-amm (panel b and d). The PTC is defined as the distance difference between the O-H and N-H in L-gln (panel e) and between the two O-H involved in L-pyro-amm (panel f).
Figure 1: Unit cell crystallographic geometries for L-gln (panel a) and L-pyro-amm (panel b), where the blue circles highlight the HBs and SHBs, respectively. Histograms of BD (panels c and d) and PTC (panels e and f) obtained from an ab-initio ground state dynamics with PBC using DFT (blue dashed line) and DFTB (red solid lines)
@@ 4.2 Absorption Spectra in Molecular Crystals
With the ground-state configurations obtained from the ab-initio MD, we next turn to extracting the absorption spectra. Figure 2 shows the absorption spectra for L-gln and L-pyro-amm obtained experimentally12 (left panel) as well as those calculated with TD-DFTB (right panel). As one can observe, an important difference between the two systems is the extent of the red-tail/edge excitation that is significantly more pronounced in the case of L-pyro-amm. In the case of L-gln, the absorption peaks are below the 250 nm and outside the window of experimentally detectable wavelengths. On other hand for L-pyro-amm, there is a very pronounced red-edge excitation tail ranging from 250-400 nm.
Figure 2: Absorption UV-Vis Spectra for L-Glutamine and L-Pyro-Ammonium. Left panel shows the experimental data, extracted from a previous study in our group12. Right panel shows the calculated average spectra using TD-DFTB. Red solid lines correspond to L-pyro-amm system while blue dashed lines for L-gln.
The right panel of Figure 2 depicts the average calculated absorption spectra obtained for both systems using 1000 averaged snapshots each, at the level of TD-DFTB (full-range spectra are shown in Figure S2 in the Supplementary Information). Interestingly, we observe that TD-DFTB reproduces the correct experimental trends comparing L-gln and L-pyro-amm. Analysing the transitions involved in the UV-Vis energy range (330-370 nm) we found that the main molecular orbitals that explain these transitions involve significant contributions coming from the carbonyl bonds and the SHB (see Figure S3 in the Supplementary Information). It is also worth noting that this prediction is also fully consistent with the results that were previously obtained at TD-DFT level in our previous work12.
@@ 4.3 Excited State TSH Simulations in Gas phase
Figure 3: Average populations of the electronic states obtained from TSH trajectories for L-gln (left panel) and L-pyro-amm (right panel) dimer systems in vacuum. Results using TD-DFT are depicted as dashed lines while those obtained with TD-DFTB in solid lines. The evolution of $S_{1}$ and $S_{0}$ are denoted in red and blue lines, respectively.
@@
In our previous studies employing TD-DFT, we observed that the transition from $S_{1}\to S_{0}$ occurs via a conical intersection (CoIn) where the essential vibrational modes coupled to the process involves the extension of the carbonyl bond, proton transfer along the SHB, and finally, the deplanarization of the amide group. In this context, it is recognized that single-determinant methods such as TD-DFT (or TD-DFTB) present an inaccurate description of the $S_{1}/S_{0}$ CoIn. However, in our prior work (see Figure 5 in the SI in the reference16), we validated the CoIn identified with TD-DFT using the CASPT2 method. We found very good agreement between both theory levels.
Examining each of the modes mentioned before in the ground and excited state as well as in the CoIn, a very specific structural signature becomes apparent. Figure 4 depicts the histograms for each of these modes obtained from TSH simulations employing both TD-DFT (top) and TD-DFTB (bottom) theory levels for L-gln dimer. The histograms were computed using data from all TSH trajectories. At each time step, we identified the current potential energy surface ($S_{1}$ or $S_{0}$), or if the system transitioned to the other surface (CoIn).
@@
Figure 4: Structural features characterizing ground (blue dashed line) and excited states (red solid line) as well as CoIn (green dashed-dot line) crossing for L-gln system. Molecular structures showing the parameters to describe the carbonyl stretching (panel a), the proton transfer (panel d) and the deplanarization (panel g). Probability Distribution Function (PDF) for the CO distance (panels b and c), PTC (panels e and f) and Deplanarization modes (panels h and i) for L-gln dimers in vacuum obtained from TSH trajectories using TD-DFT (middle panels) and TD-DFTB (lower panels). All PDFs were normalized independently.
@@
The comparison between TD-DFTB and TD-DFT reveals excellent agreement both in terms of the overall trends observed but also in the specific details of how the modes change during non-radiative decay. The variation observed in the maximum peak intensities of the PDFs for both $S_{1}$ and $S_{0}$ across all depicted modes in Figure 4 arises from the significant displacement between the minima on both potential energy surfaces. This displacement is a consequence of the substantial distortions experienced by the system upon excitation, a feature effectively captured by the DFTB approach. In particular, we observe that an additional stretching for CO respect to $S_{1}$ distribution is necessary for the system to reach the CoIn (green dashed-dot lines in panel b and c). Similar behavior is also observed for the PTC (panels e and f). The primary difference arises in relation to the deplanarization mode. In TD-DFT (panel h), we observe a higher degree of deplanarization at the CoIn compared to the distribution of this mode in the $S_{0}$ state. In contrast, although the distributions obtained with TD-DFTB reflect the same trends (panel i), the deplanarization mode appears to be much more constrained compared to our previous findings. Despite the differences, it should be underscored that, based in our previous work16, the deplanarization mode is less critical for the photophysics of both L-pyro-amm and L-gln.
@@
We next turn to comparing the behavior of the non-radiative decay modes that are important for L-pyro-amm. Since this system is fluorescent, extracting the decay mechanisms that emerge from the excited state simulations is particularly challenging due to the fact that there are fewer statistics of reactive events. In our previous study16 for example, out of a total of 100 trajectories, 4 displayed non-radiative decay from the excited to the ground-state.
@@
Figure 5: Structural features characterizing ground (blue dashed line) and excited states (red solid line) as well as CoIn (green dashed-dot line) crossing for L-pyro-amm system. Molecular structures showing the parameters to describe the carbonyl stretching and the proton transfer are in panel a and d, respectively. Probability Distribution Function (PDF) for the CO distance (panels b and c), PTC (panels e and f) for L-pyro-amm dimers in vacuum obtained from TSH trajectories using TD-DFT (middle panels) and TD-DFTB (lower panels). All PDFs were normalized independently.
@@
In Figure 5 we show the most relevant features, namely the carbonyl stretching and the proton transfer modes obtained from the TSH simulations at TD-DFT and TD-DFTB levels. Specifically, we are able to reproduce the significant overlap in the distribution of the two modes between the $S_{1}$ and $S_{0}$ electronic states (red solid and blue dashed lines). These findings indicate that the energy minimum in the potential energy surfaces for each of these modes do not exhibit significant displacement, suggesting that these modes are restricted in L-pyro-amm. This feature is also fully consistent with typical aromatic fluorescent systems69, where the restriction arises from the high electron delocalization in the planar arrangement of these compounds. In the case of non-aromatic systems, like the ones we are tackling, the strong interactions due to SHBs induce this restriction on different modes.
@@
It is important to highlight when comparing the top and bottom panels of Figure 5, that due to the reduced computational complexity, we are able to run significantly more TSH trajectories compared to our previous TDDFT simulations (600 vs 100). DFTB thus permits the observation of many more non-radiative decays yielding statistics which allow the construction of histograms of various modes near the CoIn, for instance the green distribution in the case of TD-DFTB represents 122 decays (panels c and f) and the green vertical lines for TD-DFT represents only 4 decays (panels b and e). The examination of the distribution of these modes at the CoIn reveals that the mechanism observed in L-pyro-amm is similar to that observed in the L-gln system. Specifically, we find with the DFTB approach that the CO stretching mode undergoes two distinct signatures at the CoIn. One exhibits a more subtle stretching in comparison to the distribution of $S_{1}$, while the other shows a more significant stretching of approximately $1.6$(green dashed-dot line in panel c in Figure 5). Our previous TD-DFT simulations sample a few events between these two peaks.
@@
The effect of having more statistics with DFTB is observed more clearly in the PTC distributions (panels e and f) which display more structure compared to those obtained with TD-DFT. For example, we can capture the double-well character of the PTC (blue line in panel f) with TD-DFTB, while with TD-DFT we observe one main peak along with a shoulder (blue dashed line panel e). A distinction emerges when comparing the behavior of proton transfer between the excited and ground states in TD-DFTB. Notably, in the excited state, there is a significant reduction in proton transfer events, evidenced by the presence two peaks located at ±$\pm 0.5$ with different intensity in the PDF (depicted by the red line in panel f). A similar trend is observed in TD-DFT, albeit with less intensity. In this case, a narrow distribution is noted in $S_{1}$ compared to the ground state, with the hydrogen positioned equidistant from the two oxygen atoms involved in the SHB (illustrated by the red line in panel e). Additionally, we observe that the SHB experiences a deformation at the CoIn (green dashed-dot line in panel f in Figure 5), transforming into a transient normal HB. This is evident from the peaks at higher values of approximately ±$\pm 2.5$, which is also present in TD-DFT (green dashed-dot lines in panel e). Although we cannot rule out the possibility that some of these differences arise from the underlying potential energy surface along various vibrational modes, the limited statistics with TD-DFT are more likely to be the source of the differences we observe. However, we can see that the mechanism predicted by TD-DFTB is in excellent agreement with our previous studies12, 16.
@@ 4.4 Optical Properties in Realistic Biological Environments
The preceding results provide evidence that DFTB is a very viable approach for studying the non-radiative decay mechanisms in organic molecular crystals implicated in non-aromatic fluorescence. As previously mentioned, to simulate a more realistic system beyond vacuum dimers, we adopted the QM/MM framework to account for environmental effects. Consequently, in the following section we utilize TD-DFTB to investigate the optical properties of L-gln and L-pyro-amm within this approach.
4.4.1 L-Pyroglutamine-Ammonium
@@
The experimental data used to compare the optical properties in this system was the excitation and emission spectra, as reported in ref 12. It can be observed that the maximum of the experimental excitation spectrum is around 350 nm (red solid lines in the right panel in Figure 7), which corresponds to the low lying transitions in the absorption spectrum in L-pyro-amm crystal (red solid line in the left panel of Figure 2). Therefore, to allow for a one-to-one comparison, we determined the excitation spectra instead of the absorption, i.e., including only the lower energy transitions in our calculations. We first decided to compute the excitation spectra using only the first transition ($S_{0}\to S_{1}$). Calculated and experimental results are shown in Figure 7.
@@
Figure 7: Excitation (solid lines) and Fluorescence (dashed lines) spectra for L-pyro-amm. Left panel shows the average spectra obtained from QM/MM simulations using the dimer (red line) and the unit cell (blue line) in the QM region. Right panel shows the experimental spectra (extracted from reference 12, red line) and the average calculated from the QM/MM simulations for the unit cell in the QM region, where the calculated spectra were shifted to match the maximum in the experimental fluorescence spectra. All the spectra were normalized independently.
@@
The left panel in Figure 7 shows the results for both QM systems in L-pyro-amm in the QM/MM simulation. We can observe that both systems show similar excitation maximum (approx. 215 nm) and emission (approx. 290 nm). When we compared these values with the experimental data (red lines in the right panel) we observe notable differences. This finding is a well known fact that TD-DFT and therefore in its approximate variant TD-DFTB, tend to exhibit a notable shift in the absolute magnitude of the transition energies when comparing with experimental measurements70, 71. An additional factor contributing to the discrepancy between our computational outcomes and experimental findings lies in the inherent limitations of the QM/MM approach in representing the environmental effects. As demonstrated in Section 4.2, our methodology achieves outstanding agreement with experimental results when employing a sampling coming from ab-initio ground state simulations with periodic boundary conditions (see Figure 2). Regrettably, this level of concordance is not replicated in our QM/MM simulations. To illustrate the impact of the MM environment, we conducted excited state calculations for the unit cell system using the same conformations employed for calculating the excitation spectra (blue solid line in the left panel of Figure 7), with all MM atoms removed. The results are presented in Figure S4 in the Supplementary Information. We successfully reproduced the calculated absorption spectra depicted in Figure 2 (red solid line in the right panel), thereby demonstrating the significant blue-shift effect induced by the MM environment. However, as we will elucidate in the following, these disparities persist not only in the ground state but also in excited state simulations. Consequently, focusing on energy differences becomes a more pertinent and meaningful avenue for comparison. We have thus determined the Stokes shift, defined as the difference between the maximum of the emission band and the absorption or excitation spectra the results of which are summarized in Table 1. Notably, we find remarkable agreement between the experimental results and the two model systems for the Stokes shift. It is also interesting to observe that the dimer system appears to fully capture the energetics associated with the fluorescence Stokes shift which reinforces the use of the dimer model systems in disentangling the non-radiative decay mechanisms.
@@
Table 1: Stokes shift for the different models in the QM/MM simulations
  • $a$ FWHM denotes the Full Width at Half Maximum for the fluorescence spectra.
@@
The right panel of Figure 7 presents the excitation and fluorescence spectra for both experimental and calculated using the unit cell of L-pyro-amm in the QM/MM simulations. The calculated spectra were adjusted to align with the peak of the emission spectra. Impressively, our calculations are in rather modest agreement with the experimental data. The deviation of 10 nm in the Stokes shift and also the slight shoulders at 320 nm present in the experimental excitation spectra (red solid line) which are absent in the emission (red dashed line) are an indication that the excitation spectra involves more than only one transition from the ground state. Consequently, we computed the excitation spectra by incorporating an increasing number of transitions to accommodate this behavior. We have noted that incorporating additional states leads to a slight improvement in the agreement of the Stokes shift with the experimental data, up to the inclusion of 10 states. However, beyond this point, the results show poor agreement, and the full width at half maximum (FWHM) is also not accurately reproduced in all cases (see Figure S5 and Table S1 in the Supplementary Information).